OpenStudy (anonymous):

A bag of contains 8 pink marbles, 5 green marbles and 10 brown marbles. What is the chance of drawing a pink marble? If a pink marble is drawn then placed back into the bag, and a second marble is drawn, what is the probability of drawing a green marble? Give solutions exactly in reduced fraction form, separated by a comma.

OpenStudy (amistre64):

hitting the enter key would help make this readable: A bag of contains: 8 pink marbles, 5 green marbles 10 brown marbles. What is the chance of drawing a pink marble? If a pink marble is drawn then placed back into the bag, and a second marble is drawn, what is the probability of drawing a green marble? Give solutions exactly in reduced fraction form, separated by a comma.

OpenStudy (amistre64):

23 chances in all; and 8 of those chances are pink P(pink) = 8/23

OpenStudy (amistre64):

P(pink and green) = P(pink) * P(green)

OpenStudy (amistre64):

\[\frac{8}{23}*\frac{5}{23}\]

OpenStudy (amistre64):

40 ----- 529